# BianQue: Balancing the Questioning and Suggestion Ability of Health LLMs with Multi-turn Health Conversations Polished by ChatGPT

### v1

@@ 1 Introduction
Recently, Large language models (LLMs), e.g. ChatGPT (chatgpt), LLaMA (touvron2023llama), ChatGLM (zeng2023glm-130b), have been extensively applied in various fields. Through high-quality instruction fine-tuning and reinforcement learning based on human feedback (RLHF) (instructgpt), LLMs already have possessed stunning language comprehension, generation, and knowledge reasoning abilities. Overall, users are amazed by the excellent suggestion ability of LLMs.
@@
At present, research on LLMs in the health field mainly focuses on evaluating the performance of existing models, constructing suitable datasets, and fine-tuning instructions. singhal2022large proposed a medical Q&A benchmark MultiMedQA for evaluating the clinical knowledge QA abilities of LLMs. li2023chatdoctor constructed a real doctor-patient dialogue dataset HealthCareMagic-100k, and used it to fine-tune the ChatDoctor based on LLaMA. Similar health LLMs have been released one after another, e.g. BenTsao (本草) (wang2023huatuo), ChatGLM-6B-Med (wang2023huatuo), DoctorGLM (xiong2023doctorglm), MedAlpaca (han2023medalpaca), ClinicalGPT (wang2023clinicalgpt) and etc. These models are basically based on the assumption that “users can clearly describe their problems or situations”. Therefore, during the model construction phase, the questioning ability of the model was not considered. Although these models have achieved well performance in the field of medical QA, they do not have the ability to ask users questions.
@@ 2.1 BianQueCorpus: Balancing Questioning and Suggestion
In the field of Chinese health conversational AI, there are already some multi-turn conversation datasets, e.g. MedDialog-CN (he2020meddialog), IMCS-V2 (10.1093/bioinformatics/btac817), CHIP-MDCFNPC (zhang-etal-2022-cblue), MedDG (zhang-etal-2022-cblue). However, these conversations are often crawled from internet consultation platforms, such as 好大夫 [^5] https://www.haodf.com/. These datasets are often mixed with a large amount of noise, such as missing content, missing images, reward information, privacy content, incomplete JSON content, website link, website tips, voice recording, text automatically replied by the system, etc. We first collected real-world multi-turn health conversations through data outsourcing services. Then, we performed a two-stage data optimization process: (i) We constructed a data automatic cleaning process based on regularized expression to improve the quality of existing conversation datasets. (ii) We designed a polishing prompt (see Figure 4) and use ChatGPT to polish the doctors’ suggestion of multi-turn conversations, because doctors often respond very briefly through internet platforms, lacking detailed analysis and suggestions. The whole construction process of BianQueCorpus is presented in Figure 3. We ultimately obtained a multi-turn health conversation dataset consisting of 2,437,190 samples, in which the questions accounted for 46.2% among the doctors’ answers.
@@ 2.2 BianQue Model
We chose the ChatGLM-6B (du-etal-2022-glm; zeng2023glm-130b) as the base LLM architecture to construct the BianQue, since it is open source and has excellent Chinese understanding and generation performance. The input of model is defined as:
@@ 3.1 Baselines and Benchmarks
We select ChatGLM-6B [^6] https://github.com/THUDM/ChatGLM-6B (zeng2023glm-130b), ChatGPT (gpt-3.5-turbo) (chatgpt), and DoctorGLM (xiong2023doctorglm) as the baseline models. Comparative experiments were conducted on test set of MedDialog-CN, IMCS-V2, CHIP-MDCFNPC and MedDG respectively, since they are multi-turn conversation datasets that have both suggestions and questions in targets.
@@ 3.3 Results and Analysis
Following the ClinicalGPT (wang2023clinicalgpt), We evaluated BianQue and other models with the metrics: BLEU-1/2/3/4 (papineni-etal-2002-bleu) and ROUGE-1/2/L (lin-2004-rouge). In addition, we define a new metric to measure the model’s Proactive Questioning ability (PQA):

### v2

@@ 1 Introduction
Recently, Large language models (LLMs), e.g. ChatGPT (OpenAI, 2022), LLaMA (Touvron et al., 2023), ChatGLM (Zeng et al., 2023), have been extensively applied in various fields. Through high-quality instruction fine-tuning and reinforcement learning based on human feedback (RLHF) (Ouyang et al., 2022), LLMs already have possessed stunning language comprehension, generation, and knowledge reasoning abilities. Overall, users are amazed by the excellent suggestion ability of LLMs.
@@
At present, research on LLMs in the health field mainly focuses on evaluating the performance of existing models, constructing suitable datasets, and fine-tuning instructions. Singhal et al. (2022) proposed a medical Q&A benchmark MultiMedQA for evaluating the clinical knowledge QA abilities of LLMs. Li et al. (2023) constructed a real doctor-patient dialogue dataset HealthCareMagic-100k, and used it to fine-tune the ChatDoctor based on LLaMA. Similar health LLMs have been released one after another, e.g. BenTsao (本草) (Wang et al., 2023b), ChatGLM-6B-Med (Wang et al., 2023b), DoctorGLM (Xiong et al., 2023), MedAlpaca (Han et al., 2023), ClinicalGPT (Wang et al., 2023a) and etc. These models are basically based on the assumption that “users can clearly describe their problems or situations”. Therefore, during the model construction phase, the questioning ability of the model was not considered. Although these models have achieved well performance in the field of medical QA, they do not have the ability to ask users questions.
@@ 2.1 BianQueCorpus: Balancing Questioning and Suggestion
In the field of Chinese health conversational AI, there are already some multi-turn conversation datasets, e.g. MedDialog-CN (He et al., 2020), IMCS-V2 (Chen et al., 2022), CHIP-MDCFNPC (Zhang et al., 2022), MedDG (Zhang et al., 2022). However, these conversations are often crawled from internet consultation platforms, such as 好大夫 [^5] https://www.haodf.com/. These datasets are often mixed with a large amount of noise, such as missing content, missing images, reward information, privacy content, incomplete JSON content, website link, website tips, voice recording, text automatically replied by the system, etc. We first collected real-world multi-turn health conversations through data outsourcing services. Then, we performed a two-stage data optimization process: (i) We constructed a data automatic cleaning process based on regularized expression to improve the quality of existing conversation datasets. (ii) We designed a polishing prompt (see Figure 4) and use ChatGPT to polish the doctors’ suggestion of multi-turn conversations, because doctors often respond very briefly through internet platforms, lacking detailed analysis and suggestions. The whole construction process of BianQueCorpus is presented in Figure 3. We ultimately obtained a multi-turn health conversation dataset consisting of 2,437,190 samples, in which the questions accounted for 46.2% among the doctors’ answers.
@@ 2.2 BianQue Model
We chose the ChatGLM-6B (Du et al., 2022; Zeng et al., 2023) as the base LLM architecture to construct the BianQue, since it is open source and has excellent Chinese understanding and generation performance. The input of model is defined as:
@@ 3.1 Baselines and Benchmarks
We select ChatGLM-6B [^6] https://github.com/THUDM/ChatGLM-6B (Zeng et al., 2023), ChatGPT (gpt-3.5-turbo) (OpenAI, 2022), and DoctorGLM (Xiong et al., 2023) as the baseline models. Comparative experiments were conducted on test set of MedDialog-CN, IMCS-V2, CHIP-MDCFNPC and MedDG respectively, since they are multi-turn conversation datasets that have both suggestions and questions in targets.
@@ 3.3 Results and Analysis
Following the ClinicalGPT (Wang et al., 2023a), We evaluated BianQue and other models with the metrics: BLEU-1/2/3/4 (Papineni et al., 2002) and ROUGE-1/2/L (Lin, 2004). In addition, we define a new metric to measure the model’s Proactive Questioning ability (PQA):
@@ 4 Conclusion and Future Work
Acknowledgements
This work was supported by the Science and Technology Project of Guangzhou (202103010002), the Natural Science Foundation of Guangdong Province (2022A1515011588), the National Key R&D Program of China (2022YFB4500600), the Science and Technology Project of Guangdong (2022B0101010003), the National Natural Science Foundation of China under Grant U1801262 and Guangdong Provincial Key Laboratory of Human Digital Twin (2022B1212010004).